# CT-UIO: Continuous-Time UWB-Inertial-Odometer Localization Using Non-Uniform B-spline with Fewer Anchors

### v1

@@ III-C3 IMU Factor
where $\varepsilon(\cdot)$ represents the predicting function. $\boldsymbol{\omega}_{I}(t)$ and $\textbf{a}_{I}(t)$ can be obtained from the derivatives of the continuous-time B-splines in Eq. (20), respectively.
@@
where $(\cdot)^{v}$ maps the skew-symmetric matrix to the corresponding 3D vector, $\ddot{\mathbf{p}}_{I}(t)$ represents the translational component submatrices of the second order derivatives, ${}_{I}^{W}\dot{\mathbf{R}}(t)$ is the rotational component submatrices of the first order derivatives. The continuous-time trajectory of IMU in the world frame is denoted as
@@
III-C4 Odometer Factor
We assume all intrinsic and extrinsic parameters in UWB/IMU/odometer are pre-calibrated [35, 36], so we can handily get odometer trajectory ${}_{O}^{W}\mathbf{T}(t)$. The odometer factor can be defined as
$\mathbf{e}_{t}^{odom}=\left\|\mathbf{x}_{t}-\psi\left(\mathbf{x}_{t-1},\boldsymbol{\omega}_{o}(t),\mathbf{v}_{o}(t),\eta^{o}\right)\right\|_{2},$ (38)
@@ III-C4 Odometer Factor
III-D Adaptive Sliding-Window Continuous-time Trajectory Optimization
@@
We design a continuous-time factor graph optimization for robot trajectory estimation based on the association between UWB/IMU/odometer data stream. Non-uniform B-splines are employed to obtain the optimal trajectory estimation continuously over time. To keep the computation tractable for online performance, an adaptive sliding window is designed to maintain a limited number of control points. The window length is adaptively updated based on the control point density as follows:
@@
$\Delta L=\left\{\begin{array}[]{lc}L_{\min}&\alpha\geq\lambda_{1},\\ \beta\frac{\left(L_{\max}-L_{\min}\right)\left(\lambda_{1}-\alpha\right)}{\lambda_{1}-\lambda_{0}}&\lambda_{1}<\alpha<\lambda_{2},\\ L_{\max}&\alpha\leq\lambda_{0},\end{array}\right.\$ (41)
where $\alpha$ is the control point density
$\alpha=\frac{n}{\triangle T}.$ (42)
@@
$\triangle T$ is the given time interval, $n$ is the number of control points in the trajectory segment in $[t,t+\triangle T]$. $\lambda_{0}$ and $\lambda_{1}$ denote the upper and lower thresholds of $\alpha$, $L_{min}$ and $L_{max}$ denote the minimum and maximum length values of the sliding window, respectively. The setting of $L_{min}$ and $L_{max}$ are adjusted based on the intrinsic correlations of historical data and the the system’s available resources. $\beta$ is a constant related to the measurement frequency.
@@
The factor graph optimization is solved consists of UWB ranging factor, IMU factor, odometer factor, the optimal state $\hat{\mathbf{X}}$ over an adaptive window of recent $\tau$ control points can be solved as follows:
@@
$\begin{split}\hat{\mathbf{X}}=\arg\min\sum_{k=t-\tau}^{t}\Bigg{(}&\left\|\mathbf{e}_{k}^{\text{Range}}\right\|_{\Sigma_{k}^{\text{Range}}}^{2}+\left\|\mathbf{e}_{k}^{\text{IMU}}\right\|_{\Sigma_{k}^{\text{IMU}}}^{2},\\ &+\left\|\mathbf{e}_{k}^{\text{Odom}}\right\|_{\Sigma_{k}^{\text{odom}}}^{2}+\left\|\mathbf{e}_{k}^{\text{Prior}}\right\|_{\Sigma_{k}^{\text{prior}}}^{2}\Bigg{)},\end{split}$ (43)
$\mathbf{X}^{\text{prior }}=\left\{\mathbf{x}_{t-\tau-\kappa},\ldots,\mathbf{x}_{t-\tau},\Sigma_{t}^{\text{Prior }}\right\},$ (44)
@@
where $\sum_{k}^{\text{Range}}$, $\sum_{k}^{\text{IMU}}$, $\sum_{k}^{\text{Odom}}$, $\sum_{k}^{\text{Prior}}$ is the covariance matrix associated with UWB ranging, IMU, odometer measurements, and prior factor from marginalization, respectively. $\textbf{X}^{\text{prior}}$ represents the set of control points shared by adjacent adaptive sliding windows. We maintain a temporal sliding window within an adaptive length, where the active control points within the current window are optimized and updated. As the adaptive window slides forward, old control points $\textbf{x}_{t-\tau-\kappa}$ are removed from the sliding window. However, they are still involved in constructing residuals within the current window to ensure continuance in motion estimation.
@@
Figure 4: The experimental configurations in the corridor and exhibition hall environments. The red circles represent the positions of the UWB anchor.
IV Experiments
IV-A Experimental Setup
@@
Considering that there are no publicly available datasets for the UWB-Inertial-odometer system, we validated the effectiveness of the proposed CT-UIO on self-collected datasets. Experimental campaigns are conducted in the corridor and exhibition hall scenes with different numbers of UWB anchors, as shown in Fig. 4. The UWB tag, namely, Nooploop LinkTrack, IMU, and odometer are integrated into the Turtlebot3 Waffle Pi, as shown in Fig. 1. The maximum translational velocity and rotational velocity of the robot are $0.26m/s$ and $1.82rad/s$, respectively. The calibration of spatial position between these sensors is predefined. The measurement rate for each UWB anchor is set to $32Hz$, while odometer and IMU measurements are recorded at $28Hz$ and $127Hz$, respectively. UWB anchors are mounted on the tripods at a preset height, and the position of UWB anchors deployed in the environment has been calibrated.
The proposed CT-UIO is compared with three state-of-the-art UIO methods in the same experimental environments. These include discrete-time methods (Optimization+Filtering [26] and Optimization+Trust Region [27]), uniform B-spline-based continuous-time method (SFUISE) [17]. The robot trajectories, generated by the LaMa framework [^1] https://github.com/iris-ua/iris_lama through the fusion of LIDAR and IMU data, are regarded as the ground truth. All methods are designed with C++ and operated on the robot with the CPU of ARMv8 (NVIDIA JetsonTX2) in Melodic ROS (Robot Operating System).
@@
(a) A-S1
(b) A-S2
(c) A-J1
(d) A-J2
(e) A-H1
(f) A-H2
IV-A1 Corridor Dataset
The proposed algorithm is first evaluated in the corridor environment, where four UWB anchors are set within a $16.2m\times 15.6m$ area. These anchors are placed at the corners of the area, with coordinates $(0.88,2.99)m$, $(3.36,13.64)m$, $(15.01,-0.19)m$, and $(17.03,10.65)m$, respectively. Due to physical barriers, such as walls or metal structures, between UWB anchors, only one or two UWB anchors are available under this non-fully observable condition. The robot’s motion mode is varied by adjusting its linear acceleration. A total of six sequences are collected, consisting of three types of motion modes, that is slow, fast, and hybrid, as shown in Table I.
TABLE I: Description of the sequences of the corridor dataset.
Seq Duration (second) Length (m) Translational velocity max/min (m/s) Description
@@ IV-A1 Corridor Dataset
TABLE II: Performance comparison in APE (RMSE, Meter) of different methods on the corridor dataset. The best results are marked in bold.
Seq Optimization+ Filtering[26] Optimization+ Trust Region[27] SFUISE [17] Ours
A-S1 0.756 0.550 0.210 0.208
A-S2 0.605 0.520 0.262 0.242
A-J1 0.248 0.575 0.229 0.181
A-J2 0.382 0.593 0.208 0.153
A-H1 0.447 0.658 0.487 0.403
A-H2 0.548 0.671 0.448 0.416
@@
The results are summarized in Table II, showing that our CT-UIO achieves the best accuracy when compared with the state-of-the-art methods. Due to the lack of sufficient available UWB anchors in the corridor environment, leading to fewer constraints in the pose graph optimization process and reduced localization accuracy for Optimization+Filtering [26] and Optimization+Trust Region [27] methods. In contrast, our CT-UIO introduces virtual UWB anchors to ensure full observability in the few-anchor localization system. This allows for the incorporation of multiple constraints into the factor graph optimization, resulting in more accurate localization estimates.
@@
In sequences such as A-S1 and A-S2, the improvement in APE is relatively minor when compared with the uniform B-spline-base method SFUISE. In these sequences, where the robot’s movement is relatively slow, the proposed adaptive knot span adjustment does not necessarily lead to a significant improvement in performance. However, in sequences involving fast linear motion, such as A-J1 and A-J2, the errors in SFUISE increase significantly. In contrast, CT-UIO outperforms state-of-the-art methods in terms of the overall performance of pose interpolation. Across hybrid sequences, CT-UIO can improve localization accuracy significantly by $38.75\%$ in APE. This improvement is attributed to the adaptive knot span adjustment strategy, which allows dynamic adjustment of the number of control points by either lowering or increasing the knot span to accommodate varying motion speeds. Fig. 5 shows the overview of trajectories in the corridor dataset, our method achieves consistent moving trajectories in most sequences. Fig. 6 shows the illustration of the proposed adaptive knot span adjustment strategy in sequence A-H2.
Figure 6: The distribution of control points in sequence A-H2.
IV-B Exhibition Hall Dataset
@@
We also evaluate CT-UIO and competing state-of-the-art methods on the exhibition hall dataset. In this case, two UWB anchors are placed at positions $(4.52,10.68)m$, $(-1.71,-2.29)m$ to create the non-fully observable condition within an area of $20.0m\times 11.6m$ and there is no occlusion between UWB anchors. To evaluate the effectiveness of the methods under aggressive turns, we provide the corresponding range of rotational velocity readings, which indicate instances of aggressive motion. A total of six sequences are detailed in Table III.
TABLE III: Description of the sequences of the exhibition hall dataset.
Seq Duration (second) Length (m) Rotational velocity max/min (rad/s) Description
@@
  • Fast sequences: the robot moves at large rotational velocity with intense speed variations, following a wave path.

### v2

@@ IV-C2 IMU Factor
where $\xi(\cdot)$ represents the pre-integration of IMU measurements. $\boldsymbol{\omega}_{I}(t)$ and $\textbf{a}_{I}(t)$ can be obtained from the derivatives of the continuous-time B-splines in Eq. (30), respectively.
@@
where $(\cdot)^{v}$ maps the skew-symmetric matrix to the corresponding 3D vector, $\ddot{\mathbf{p}}_{I}(t)$ represents the translational component submatrices of the second order derivatives, $_{I}^{W}\dot{\mathbf{R}}(t)$ is the rotational component submatrices of the first order derivatives. The continuous-time trajectory of the IMU in the world frame is denoted as:
@@
IV-C3 Odometer Factor
We assume all intrinsic and extrinsic parameters in UWB/IMU/odometer are pre-calibrated [bai2022graph, LI2024115186], so we can handily get odometer trajectory $_{O}^{W}\mathbf{T}(t)$. The odometer factor can be defined as
$\mathbf{e}_{t}^{\text{odom}}=\left\|\mathbf{x}_{t}-\psi\left(\mathbf{x}_{t-1},\boldsymbol{\omega}_{o}(t),\mathbf{v}_{o}(t),\eta^{o}\right)\right\|_{2},$ (38)
@@ IV-C3 Odometer Factor
IV-C4 Adaptive Sliding Window Factor Graph Optimization
@@
We design a continuous-time factor graph optimization for robot trajectory estimation based on the association between UWB/IMU/odometer data streams. Non-uniform B-splines are employed to obtain the optimal trajectory estimation continuously over time. To keep the computation tractable for online performance, an adaptive sliding window is designed to maintain a limited number of control points. The window length is adaptively updated based on the control point density as follows:
@@
$\Delta L=\left\{\begin{array}[]{lc}L_{\min}&\alpha_{cp}\geq\lambda_{1},\\ \left\lfloor\beta\frac{\left(L_{\max}-L_{\min}\right)\left(\lambda_{1}-\alpha_{cp}\right)}{\lambda_{1}-\lambda_{0}}\right\rfloor&\lambda_{0}<\alpha_{cp}<\lambda_{1},\\ L_{\max}&\alpha_{cp}\leq\lambda_{0}.\end{array}\right$.\\ (41)
Here, $\alpha_{cp}$ is the control point density
$\alpha_{cp}=\frac{n}{\triangle T}.$ (42)
@@
$\triangle T$ is the given time interval, $n$ is the number of control points in the trajectory segment in $t,t+\triangle T]$. $\left\lfloor\cdot\right\rfloor$ is the floor function, $\lambda_{0}$ and $\lambda_{1}$ denote the upper and lower thresholds of $\alpha_{cp}$, $L_{\text{min}}$ and $L_{\text{max}}$ denote the minimum and maximum length values of the sliding window, respectively. The settings of $L_{\text{min}}$ and $L_{\text{max}}$ are adjusted based on the intrinsic correlations of historical data and the system’s available resources. $\beta$ is a constant related to the IMU measurement frequency.
@@
The factor graph optimization is solved, consisting of UWB ranging factor, IMU factor, odometer factor, and the optimal state $\hat{\mathbf{X}}$ over an adaptive window $t-\tau,t]$ of recent $\Delta L$ control points can be solved as follows:
@@
$\begin{split}\hat{\mathbf{X}}=\arg\min\sum_{k=t-\tau}^{t}\Bigg(&\left\|\mathbf{e}_{k}^{r}\right\|_{\Sigma_{k}^{r}}^{2}+\left\|\mathbf{e}_{k}^{\text{IMU}}\right\|_{\Sigma_{k}^{\text{IMU}}}^{2},\\ &+\left\|\mathbf{e}_{k}^{\text{odom}}\right\|_{\Sigma_{k}^{\text{odom}}}^{2}+\left\|\mathbf{e}_{k}^{\text{Prior}}\right\|_{\Sigma_{k}^{\text{prior}}}^{2}\Bigg),\end{split}$ (43)
$\mathbf{X}^{\text{prior }}=\left\{\mathbf{x}_{t-\tau},\ldots,\mathbf{x}_{t-\tau+L_{w}}\right\},$ (44)
@@
where $\sum_{k}^{r}$, $\sum_{k}^{\text{IMU}}$, $\sum_{k}^{\text{odom}}$, $\sum_{k}^{\text{Prior}}$ is the covariance matrix associated with UWB ranging, IMU, odometer measurements, and prior factor from marginalization, respectively. $\textbf{X}^{\text{prior}}$ represents the prior factor in $t-\tau,t-\tau+L_{w}]$ shared by adjacent adaptive sliding windows. We maintain a temporal sliding window within an adaptive length, where the active control points within the current window are optimized and updated. As the adaptive window slides forward, old control points are removed from the sliding window. However, they are still involved in constructing residuals within the current window to ensure continuity in motion estimation.
@@
Figure 4: The experimental configurations in the corridor, exhibition hall, and office environments. The red circles represent the positions of the UWB anchors.
V Experiments
V-A Experimental Setup
@@
Considering that there are no publicly available datasets for the UWB-Inertial-odometer system, we validated the effectiveness of the proposed CT-UIO on self-collected datasets. Experimental campaigns are conducted in the corridor, exhibition hall, and office scenes with different numbers of UWB anchors, as shown in Fig. 4. The UWB tag, namely, Nooploop LinkTrack, IMU, and odometer are integrated into the Turtlebot3 Waffle Pi, as shown in Fig. 1. Before localization, the parameters of the UWB/IMU/odometer system have been calibrated. In LOS case, the overall UWB ranging error follows a Gaussian distribution, which can be fitted using $\mathcal{N}(-0.0552,0.0514^{2})$(unit: m). The odometer scale factor error is set to be 0.5%. The specification and noise characteristics of the IMU are summarized in Table  I.
TABLE I: Noise components in IMU
Noise
Description
Velocity
Random walk
Acc. Bias
Angular Random
walk
Gyro Bias
IMU 0.035 $m/s/\sqrt{h}$ 0.2 mgmg 1.03 $deg/h$ 8.23 $deg/h$
The measurement rate for each UWB anchor is set to $32Hz$, while odometer and IMU measurements are recorded at $28Hz$ and $127Hz$, respectively. UWB anchors are mounted on the tripods at a preset height, and the position of the UWB anchors deployed in the environment has been calibrated.
The proposed CT-UIO is compared with five state-of-the-art UIO methods in the same experimental environments. These include discrete-time methods (Optimization+Filtering [yang2023novel], Optimization+Trust Region [zhou2024optimization], GPDA [WOS:001173317800046], and VMHE [9829196]) and uniform B-spline-based continuous-time method (SFUISE [li2023continuous]). The robot trajectories, generated by the LaMa framework [^1] https://github.com/iris-ua/iris_lama through the fusion of LIDAR and IMU data, are regarded as the ground truth. All methods are designed with C++ and operated on the robot with the CPU of ARMv8 (NVIDIA JetsonTX2 @ 32 GHz, with 8 GB memory) in Melodic ROS (Robot Operating System). In all experiments, the specific parameters of CT-UIO are listed in Table II. These parameter values are determined based on the noise characteristics of the onboard sensors and the computational constraints of the robot’s embedded processor. The selection aims to balance the trade-offs between localization accuracy and computational efficiency. To further investigate the influence of dataset characteristics on the selection of $\tau_{F}$, we performed the parameter analysis separately for the three datasets with varying numbers and layouts of anchors. For each dataset, a parameter sweep of $\tau_{F}$ was conducted on the MATLAB platform using a series of simulations with various VAs’ positions. For each configuration, we evaluated the initialization success rate, the number of UWB observations, the time-to-initialization, and the waypoints obtained from the IMU/odometer fusion model. The optimal threshold $\tau_{F}$ is then determined as the mid-point of the margin of minimum eigenvalues of well-conditioned waypoints and ill-conditioned ones.
TABLE II: Values of parameters used in experiments.
Parameter Value
A constant related to the IMU measurement noise frequency $\alpha$
$\alpha_{\text{acc}}=1.22\times 10^{-4}$
$\alpha_{\text{gyro}}=3.17\times 10^{-3}$
Bounds for the $e(k)$
$\lambda_{\text{min}}=0.56$
$\lambda_{\text{max}}=6.52$
The maximum size of IMU observation statistics $\xi=50$
Threshold for the rejection domain of $E(k)$ $thr=10.85$
Outlier rejection threshold for the individual ranging measurement $\gamma=1.25$
The lower bound of the VA’s position estimation error
Corridor $\tau_{F}=25$
Exhibition Hall $\tau_{F}=44$
[…]
Bounds for the sliding window
$L_{\text{min}}=5$
$L_{\text{max}}=30$
Bounds for the control point density $\alpha_{cp}$
$\lambda_{0}=0.02$
$\lambda_{1}=0.1$
The translational-velocity-change bins
$\lambda_{v_{1}}=0.06,\lambda_{v_{2}}=0.12$
$\lambda_{v_{3}}=0.20,\lambda_{v_{4}}=0.26$
The rotational-velocity-change bins
$\theta_{\omega_{1}}=0.20,\theta_{\omega_{2}}=0.50$
$\theta_{\omega_{3}}=0.13,\theta_{\omega_{4}}=1.82$
@@
V-B Evaluation on the Corridor Dataset
((a))
((b))
((c))
((d))
((e))
((f))
Figure 5: We query the estimated trajectory from the obtained continuous-time trajectory and evaluate it against the ground-truth trajectory. The ground-truth trajectory is shown as a black dashed line, and the estimated trajectory is colored by absolute translation error when compared against the ground truth in the corridor dataset. Our proposed CT-UIO not only achieves high accuracy in trajectory estimation but also effectively fits translational velocity variations.
The proposed algorithm is first evaluated in the corridor environment, where four UWB anchors are set within a $16.2m{\times}15.6m$ area. These anchors are placed at the corners of the area, with coordinates $(0.88,2.99)m$, $(3.36,13.64)m$, (15.01,0.19)m(15.01,{-}0.19)m, and $(17.03,10.65)m$, respectively. Due to physical barriers, such as walls or metal structures, between UWB anchors, only one or two UWB anchors are available under this non-fully observable condition. The robot’s motion mode is varied by adjusting its linear acceleration. A total of six sequences are collected, consisting of three types of motion modes, that is slow, fast, and hybrid, as shown in Table III.
TABLE III: Description of the sequences of the corridor dataset.
Seq
Duration
Length
Translational velocity
max/min (m/s)
Description
@@ V-B Evaluation on the Corridor Dataset
((a))
((b))
((c))
((d))
((e))
((f))
Figure 6: CDFs of APE for different methods on the corridor dataset. The APE values of our method are more concentrated below $0.5m$ compared to the other five methods. This result further demonstrates the effectiveness of the proposed method in achieving a better localization accuracy due to its adaptive knot span adjustment strategy and VA generation mechanism.
@@
TABLE IV: Performance comparison in APE (RMSE, Meter) of different methods on the corridor dataset. The best results are marked in bold.
Seq
A-S1
A-S2 A-J1 A-J2 A-H1 A-H2
Optimization+
Filtering[yang2023novel]
0.756 0.605 0.644 0.623 0.608 0.548
Optimization+
Trust Region[zhou2024optimization]
0.550 0.520 0.575 0.593 0.658 0.671
GPDA[WOS:001173317800046] 0.617 0.338 0.523 0.605 0.537 0.590
VMHE[9829196] 0.467 0.313 0.433 0.539 0.501 0.499
SFUISE[li2023continuous] 0.210 0.262 0.229 0.208 0.487 0.448
Ours 0.208 0.242 0.181 0.153 0.403 0.416
@@ A-S2 A-J1 A-J2 A-H1 A-H2
The results are summarized in Table IV, showing that our CT-UIO achieves the best accuracy when compared with the state-of-the-art methods. Due to the lack of sufficient available UWB anchors in the corridor environment, leading to fewer constraints in the pose graph optimization process and reduced localization accuracy for Optimization+Filtering [yang2023novel], Optimization+Trust Region [zhou2024optimization] and GPDA[WOS:001173317800046] methods. Despite the additional radial velocity constraint introduced in VMHE [9829196], the formulation still suffers from ambiguity and drift when the trajectory is dominated by tangential motion. In contrast, our CT-UIO introduces VAs to ensure full observability in the few-anchor localization system. This allows for the incorporation of multiple constraints into the factor graph optimization, resulting in more accurate localization estimates.
@@
In sequences such as A-S1 and A-S2, the improvement in APE is relatively minor when compared with the uniform B-spline-based method SFUISE. In these sequences, where the robot’s movement is relatively slow, the proposed adaptive knot span adjustment does not necessarily lead to a significant improvement in performance. However, in sequences involving fast linear motion, such as A-J1 and A-J2, the errors in SFUISE increase significantly. In contrast, CT-UIO outperforms state-of-the-art methods in terms of the overall performance of pose interpolation. Across hybrid sequences, CT-UIO can improve localization accuracy significantly by $38.75\%$ in APE. This improvement is attributed to the adaptive knot span adjustment strategy, which allows dynamic adjustment of the number of control points by either lowering or increasing the knot span to accommodate varying motion speeds. Fig. 5 shows the overview of trajectories in the corridor dataset, where our method achieves consistent moving trajectories in most sequences.
The positioning errors are further evaluated using cumulative distribution functions (CDFs), as shown in Fig. 6. Fig. 7 shows the illustration of the proposed adaptive knot span adjustment strategy in sequence A-H1. For sequence A-H1, $90\%$ of the APE in our method are less than $0.515m$, while for Optimization+Filtering, Optimization+Trust Region, GPDA, VMHE, and SFUISE, the values are $0.643m$, $0.645m$, $0.590m$, $0.642m$, and $0.706m$, respectively. The results demonstrate that our method achieves high levels of positioning accuracy.
Figure 7: The distribution of control points in sequence A-H1.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 8: We align the estimated trajectory for all sequences in the exhibition hall dataset with the ground-truth trajectory at the same frequency. The ground-truth trajectory is shown as a black dashed line, and the estimated trajectory is colored by absolute translation error when compared against the ground truth. It can be observed that our CT-UIO achieves high accuracy in trajectory estimation, closely matching the ground-truth trajectories across all sequences, even in those involving fast turns.
V-C Evaluation on the Exhibition Hall Dataset
@@
We also evaluate CT-UIO and competing state-of-the-art methods on the exhibition hall dataset. In this case, two UWB anchors are placed at positions $(4.52,10.68)m$, (1.71,2.29)m({-}1.71,{-}2.29)m to create the non-fully observable condition within an area of $20.0m{\times}11.6m$, and there is no occlusion between UWB anchors. To evaluate the effectiveness of the proposed method under aggressive turns, we provide the corresponding range of rotational velocity readings, which indicate instances of aggressive motion. A total of six sequences are detailed in Table V.
TABLE V: Description of the sequences of the exhibition hall dataset.
Seq
Duration
Length
Rotational velocity
max/min (rad/s)
Description
@@
  • Fast sequences: the robot moves at a large rotational velocity with intense speed variations, following a wave path.
@@
TABLE VI: Performance comparison in APE (RMSE, Meter) of different methods on the exhibition hall dataset. The best results are marked in bold.
Seq
B-S1
B-S2 B-J1 B-J2 B-H1 B-H2
Optimization+
Filtering[yang2023novel]
0.432 0.372 0.760 0.391 0.885 0.728
Optimization+
Trust Region[zhou2024optimization]
0.393 0.405 0.588 0.562 0.543 0.413
GPDA[WOS:001173317800046] 0.319 0.372 0.523 0.460 0.537 0.409
VMHE[9829196] 0.255 0.347 0.433 0.380 0.501 0.373
SFUISE[li2023continuous] 0.042 0.127 0.243 0.168 0.203 0.358
Ours 0.040 0.122 0.219 0.098 0.150 0.311
((a))
((b))
((c))
((d))
((e))
((f))
Figure 9: CDFs of APE for different methods on the exhibition hall dataset. The results of SFUISE with uniform B-splines are closer to those of our method, but its performance holds only for slow sequences, and cannot maintain consistent in fast and hybrid sequences. In contrast, our method shows a more concentrated APE distribution across all sequences compared to the other methods.